A bird moves with a velocity of 20 metre per second in a direction making an angle of 60 degree with the Eastern line and 60 degree with vertical of word represent the velocity vector in rectangular form. 

Grade:11

1 Answers

Aditya Gupta
4 years ago
let v= ai + bj + ck
then, a^2+b^2+c^2= 20^2= 400
also, v.i= 20cos60
or a= 10
also, v.k= 20cos60
or c= 10
so, b= ± root(200)= ± 10*root2
so, velocity vector in rectangular form is
v= 10(i+j ± kroot2)
kindly approve :=)